OpenStudy (anonymous):

\[\sqrt{-6}*\sqrt{7}\] perform the calculations and place the result in standard form.

OpenStudy (anonymous):

\[[\sqrt{-6}*\sqrt{7}\]

OpenStudy (anonymous):

o it did work sorry for posting twice

OpenStudy (anonymous):

\[i\sqrt{42}\]

OpenStudy (anonymous):

awesome thanks